OpenStudy (anonymous):

x + y = 5 0.1(x) + 0.25(y) = 0.2 (5)

OpenStudy (anonymous):

multiply both the equations with 100 in the first step @snsd9

OpenStudy (anonymous):

@snsd9 are you done with this step or not.. ?? or should i write down all the steps.. ?

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

That would help me a lot.

OpenStudy (anonymous):

just try solving the question here in the comments side by side that would let you understand much better so , to start with multiply both the equations with 100 i.e. 100 ( x + y ) = 100 (5) and the other 100(0.1(x) + 0.25(y)) = 0.2 (5) (100)

OpenStudy (anonymous):

Thank you Then then solve for x and y right?

OpenStudy (anonymous):

yeah from either of the equation get the equation to find out any of the x or y and then put that value in the other equation to get the values